OpenStudy (dschneider2016):

Find the standard form of the equation of the parabola with a focus at (3, 0) and a directrix at x = -3.

OpenStudy (dschneider2016):

@tkhunny

OpenStudy (tkhunny):

This is a standard question design to see if you know where the Vertex is. Draw the Focus and Directrix. It should be very easy to find the Vertex. By Definition, it must be equidistant from both. After that, we need the magic "p". That is also very easy if we know the Focus, the Vertex, and the Directrix. What distance IS the Vertex from both? Done. Go!

OpenStudy (dschneider2016):

so would it be Vertex: (0, 0); Focus: (0, 3); Directrix: y = -3; Focal width: 12?

OpenStudy (jdoe0001):

I thought the focus was given at 3,0?

OpenStudy (dschneider2016):

oh okay! I see

OpenStudy (jdoe0001):

anyhow... keeping in mind that the directrix is the same distance to the vertex as the focus is to the vertex where would that put the vertex at?

OpenStudy (jdoe0001):

also... let us notice the focus and directrix points is that a vertical or horizontal parabola anyway?

OpenStudy (dschneider2016):

Would the answer then be y = 1/12x

OpenStudy (dschneider2016):

y = 1/12x^2

OpenStudy (jdoe0001):

well.. how did you get that?

OpenStudy (dschneider2016):

Well.. Im not sure but the vertex must be at (0,0) right?

OpenStudy (jdoe0001):

hmmm 0, 0?

OpenStudy (jdoe0001):

notice the directrix notice the focus point :)

OpenStudy (dschneider2016):

the focus point is at (3,0) ad the directix is at (-3,0)

OpenStudy (dschneider2016):

so the middle of that would be (0,0)?

OpenStudy (jdoe0001):

well, yes that is correct :) also notice, the distance from the vertex to either, the directrix or the focus point, is 3 notice something  the parabola opens towards the focus, and thus, is a horizontally opening parabola thus we'll use then the form with the "y" squared or \(\begin{array}{llll} (y-{\color{blue}{ k}})^2=4{\color{purple}{ p}}(x-{\color{brown}{ h}}) \\ \\ \end{array} \qquad \begin{array}{llll} vertex\ ({\color{brown}{ h}},{\color{blue}{ k}})\\ {\color{purple}{ p}}=\textit{distance from vertex to }\\ \qquad \textit{ focus or directrix} \end{array}\) so.. what would that give you ?:)

OpenStudy (dschneider2016):

y^2=12x

OpenStudy (jdoe0001):

\(\begin{array}{llll} (y-{\color{blue}{ 0}})^2=4{\color{purple}{ (3)}}(x-{\color{brown}{ 0}}) \\ \\ \end{array} \qquad \begin{array}{llll} vertex\ ({\color{brown}{ 0}},{\color{blue}{ 0}})\\ {\color{purple}{ p}}=\textit{distance from vertex to }\\ \qquad \textit{ focus or directrix} \end{array} \\ \quad \\ y^2=12x\implies \cfrac{y^2}{12}=x\) keep in mind that because the parabola opens to the "right", the distance "p" or 3, is positive if the parabola opened to the left, it'd be -3 then

OpenStudy (dschneider2016):

y = 1/12x^2 x = 1/12y^2 -12y=x^2 y2=6x

OpenStudy (jdoe0001):

?

OpenStudy (dschneider2016):

SO out of all my choices would you agree that the 2nd option is correct? @jdoe0001

OpenStudy (jdoe0001):

yes \(\bf \cfrac{y^2}{12}=x\iff y^2\cfrac{1}{12}=x\)

OpenStudy (dschneider2016):

thanks so much!! Could you help me with another question? @jdoe0001

OpenStudy (jdoe0001):

sure, just post anew

OpenStudy (dschneider2016):

A building has an entry the shape of a parabolic arch 96 ft high and 18 ft wide at the base, as shown below. Find an equation for the parabola if the vertex is put at the origin of the coordinate system.

OpenStudy (dschneider2016):

okay!

OpenStudy (dschneider2016):

Ill post a new one